# Emergent Symmetry and Phase Transitions on the Domain Wall of $\mathbb{Z}_{2}$ Topological Orders

### v1

@@ Abstract
Despite the fairly low symmetry of the IDWM, numerical calculations suggested that its low-energy spectrum closely resembles that of the spin-$1/2$ Heisenberg chain and is consistent with the SU(2)1 WZW theory [19]. Further numerical evidence supporting the SU(2)1 WZW effective theory is provided in the Supplemental Materials (SM)  [^1] See the Supplemental Materials for details. based on the two-interval mutual information of the ground state [26, 27]. However, the microscopic origin of the emergent SU(2)1 conformal symmetry was not clear.
@@
where $R_{l}([111],2\pi/3)=\exp\big{(}i\pi(\sigma^{x}_{l}+\sigma^{y}_{l}+\sigma^{z}_{l})/3\sqrt{3}\big{)}$ is the spin rotation around the $[111]$ axis by $32\pi/$ 3 and permutes the Pauli operators cyclically, $R_{l}^{\dagger}(\sigma^{x}_{l},\sigma^{y}_{l},\sigma^{z}_{l})R_{l}=(\sigma^{y}_{l},\sigma^{z}_{l},\sigma^{x}_{l})$. This maps the IDWM into a deformed Heisenberg chain, $\tilde{H}_{0}=U_{2}^{\dagger}U_{1}^{\dagger}H_{0}U_{1}U_{2}$,
@@
The deformed Heisenberg chain in Eq. (3) has exactly the same form as the Kitaev-Gamma chain in a rotated basis studied in Ref. [28], and has a nonsymmorphic octahedral symmetry, which we explain below. It has a three-site translation symmetry $T_{3}$ ($T_{n}$ denotes the lattice translation by $n$ sites), three global spin flip symmetries $\mathcal{R}(\hat{n},\pi)=\prod_{l}R_{l}(\hat{n},\pi)$ ($\hat{n}=\hat{x}$, $\hat{y}$, $\hat{z}$), and an antiunitary time-reversal symmetry $\mathcal{T}$. Moreover, it is also invariant under the spin-screw translation $\mathcal{R}_{T}T_{1}$ and the spin-flip reflection $\mathcal{R}_{I}I_{0}$, where $\mathcal{R}_{T}=\prod_{l}R_{l}([111],-2\pi/3)$ is a spin rotation, $\mathcal{R}_{I}=\prod_{l}R_{l}([1\bar{1}0],\pi)$ is a spin flip, and $I_{0}$ denotes the site-00-centered lattice reflection. These symmetry operators generate a nonsymmorphic group $G_{0}$, which has a normal subgroup generated by the three-site translation $\langle T_{3}\rangle\cong\mathbb{Z}$, and $G_{0}/\langle T_{3}\rangle\cong O_{h}$, where $O_{h}$ denotes the full octahedral group [28].
@@
where $\mathbf{J}$ and $\bar{\mathbf{J}}$ terms are omitted because they are conserved currents and can be eliminated by redefining the boson field $\varphi_{s}$ introduced below. This symmetric bilinear term can be simplified by an orthogonal transformation, $\mathbf{M}=V\mathbf{J}$ and $\bar{\mathbf{M}}=V\bar{\mathbf{J}}$, where $V=R(\hat{x},\arctan\sqrt{2})R(\hat{z},\pi/4)$ rotates the $[111]$ axis into the $\hat{z}$ axis, and we find
@@
Figure 3: Numerical results of the domain wall model $H_{1}$. (a) Half-chain EE $S(L/2)$ versus the chain length $L$ in the logarithmic scale in the parameter range $-1.28\leq h\leq 0$. Solid lines are the fitting to $S(L/2)=(c/6)\ln L+a$, which give an estimate of central charges listed in the legend. (b) Squared FM and AF order parameters $m_{\mathrm{FM}}^{2}$ and $m_{\mathrm{AF}}^{2}$ versus $h$ for $L=400$. All calculations are performed with the open boundary condition.
We perform density-matrix renormalization group (DMRG) calculations on the domain wall model $H_{1}=H_{0}+H_{h}$. The half-chain entanglement entropy (EE) $S(L/2)$ in the parameter range $-1.28\leq h\leq 0$ is plotted in Fig. 3 (a). It obeys the formula $S(L/2)=(c/6)\ln L+a$ [33] and gives an estimate of the central charge $c\simeq 1$ as listed in the legend, which is a hallmark of the TLL state. The squared FM and AF order parameters are estimated with the spin correlation functions, $m_{\mathrm{FM}}^{2}=(2/L)\sum_{r=0}^{L/2}\langle\sigma_{L/2-r/2}^{z}\sigma_{L/2+r/2}^{z}\rangle$ and $m_{\mathrm{AF}}^{2}=(2/L)\sum_{r=0}^{L/2}(-1)^{r}\langle\sigma_{L/2-r/2}^{z}\sigma_{L/2+r/2}^{z}\rangle$, and are plotted in Fig. 3 (b). The TLL undergoes a continuous transition at $h\simeq 0$ to the FM order, which is consistent with the RG analysis. However, we find a transition at $h\simeq-1.3$ from the TLL to the AF order, which is not captured by the RG analysis. The abrupt change of the AF order parameter indicates that this is a first-order transition. The AF order spontaneously breaks the time-reversal, the lattice translation and the $\mathbb{Z}_{2}$ self-duality symmetries, hence it is also a manifestation of the anomaly of $\mathbb{Z}_{2}$ symmetry.
@@
[…]
@@
$\mathcal{H}_{g}\simeq\frac{g\gamma}{\pi}\cos\sqrt{2\pi}\varphi_{s},$

### v2

@@ Abstract
Despite the fairly low symmetry of the IDWM, numerical calculations suggested that its low-energy spectrum closely resembles that of the spin-$1/2$ Heisenberg chain and is consistent with the SU(2)1 WZW theory [19]. Further numerical evidence supporting the SU(2)1 WZW effective theory is provided in the Supplemental Materials (SM)  [^1] See the Supplemental Materials for details, which include Refs. [19, 42, 29, 30, 43, 33, 26, 27, 44]. based on the two-interval mutual information of the ground state [26, 27]. However, the microscopic origin of the emergent SU(2)1 conformal symmetry was not clear.
@@
where $R_{l}([111],2\pi/3)=\exp\big(i\pi(\sigma^{x}_{l}+\sigma^{y}_{l}+\sigma^{z}_{l})/3\sqrt{3}\big)$ is the spin rotation around the $[111]$ axis by $2\pi/3$ and permutes the Pauli operators cyclically, $R_{l}^{\dagger}(\sigma^{x}_{l},\sigma^{y}_{l},\sigma^{z}_{l})R_{l}=(\sigma^{y}_{l},\sigma^{z}_{l},\sigma^{x}_{l})$. This maps the IDWM into a deformed Heisenberg chain, $\tilde{H}_{0}=U_{2}^{\dagger}U_{1}^{\dagger}H_{0}U_{1}U_{2}$,
@@
The deformed Heisenberg chain in Eq. (3) has exactly the same form as the Kitaev-Gamma chain in a rotated basis studied in Ref. [28], and has a nonsymmorphic octahedral symmetry, which we explain below. It has a three-site translation symmetry $T_{3}$ ($T_{n}$ denotes the lattice translation by $n$ sites), three global spin flip symmetries $\mathcal{R}(\hat{n},\pi)=\prod_{l}R_{l}(\hat{n},\pi)$ ($\hat{n}=\hat{x}$, $\hat{y}$, $\hat{z}$), and an antiunitary time-reversal symmetry $\mathcal{T}$. Moreover, it is also invariant under the spin-screw translation $\mathcal{R}_{T}T_{1}$ and the spin-flip reflection $\mathcal{R}_{I}I_{0}$, where $\mathcal{R}_{T}=\prod_{l}R_{l}([111],-2\pi/3)$ is a spin rotation, $\mathcal{R}_{I}=\prod_{l}R_{l}([1\bar{1}0],\pi)$ is a spin flip, and $I_{0}$ denotes the site-0-centered lattice reflection. These symmetry operators generate a nonsymmorphic group $G_{0}$, which has a normal subgroup generated by the three-site translation $\langle T_{3}\rangle\cong\mathbb{Z}$, and $G_{0}/\langle T_{3}\rangle\cong O_{h}$, where $O_{h}$ denotes the full octahedral group [28].
@@
where terms proportional to $\mathbf{J}$ and $\bar{\mathbf{J}}$ are omitted because they are conserved currents and can be eliminated by redefining the boson field $\varphi_{s}$ introduced below. This symmetric bilinear term can be simplified by an orthogonal transformation, $\mathbf{M}=V\mathbf{J}$ and $\bar{\mathbf{M}}=V\bar{\mathbf{J}}$, where $V=R(\hat{x},\arctan\sqrt{2})R(\hat{z},\pi/4)$ rotates the $[111]$ axis into the $\hat{z}$ axis, and we find
@@
Figure 3: Numerical results of the domain wall model $H_{1}$. (a) Half-chain $E$ $S(L/2)$ versus the chain length $L$ in the logarithmic scale in the parameter range $1.28\leq h\leq 0$. Solid lines are the fitting to $S(L/2)=(c/6)\ln L+a$, which give an estimate of central charges listed in the legend. (b) Squared FM and AF order parameters $m_{\mathrm{FM}}^{2}$ and $m_{\mathrm{AF}}^{2}$ versus $h$ for $L=400$. All calculations are performed with the open boundary condition.
We perform density-matrix renormalization group (DMRG) calculations on the domain wall model $H_{1}=H_{0}+H_{h}$. The half-chain entanglement entropy (EE) $S(L/2)$ in the parameter range $1.28\leq h\leq 0$ is plotted in Fig. 3 (a). It obeys the formula $S(L/2)=(c/6)\ln L+a$ [33] and gives an estimate of the central charge $c\simeq 1$ as listed in the legend, which is a hallmark of the TLL state. The squared FM and AF order parameters are estimated with the spin correlation functions, $m_{\mathrm{FM}}^{2}=(2/L)\sum_{r=0}^{L/2}\langle\sigma_{L/2-r/2}^{z}\sigma_{L/2+r/2}^{z}\rangle$ and $m_{\mathrm{AF}}^{2}=(2/L)\sum_{r=0}^{L/2}(-1)^{r}\langle\sigma_{L/2-r/2}^{z}\sigma_{L/2+r/2}^{z}\rangle$, and are plotted in Fig. 3 (b). The TLL undergoes a continuous transition at $h\simeq 0$ to the FM order, which is consistent with the RG analysis. However, we find a transition at $h\simeq-1.3$ from the TLL to the AF order, which is not captured by the RG analysis. The abrupt change of the AF order parameter indicates that this is a first-order transition. The AF order spontaneously breaks the time-reversal, the lattice translation and the $\mathbb{Z}_{2}$ self-duality symmetries, hence it is also a manifestation of the anomaly of $\mathbb{Z}_{2}$ symmetry.
@@
Let us first highlight several exactly solvable limits. Setting g=1g=-1, the model reduces to the transverse-field Ising chain, which is exactly solvable with the Jordan-Wigner transformation [34]. Its ground state at $h=0$ is a direct-product state. This topologically trivial phase expands the range 2<h<2-2<h<2, and undergoes a continuous transition at $h=2$ to the FM order and another continuous transition at h=2h=-2 to the AF order. Both QCPs belong to the 2D Ising universality class. On the other hand, setting $g=+1$, the model reduces to the 1D cluster Ising model, which is also exactly solvable with the Jordan-Wigner transformation [35, 36]. Its ground state at $h=0$ is a $\mathbb{Z}_{2}^{T}$-SPT phase, which expands the range 2<h<2-2<h<2, and undergoes 2D Ising transitions to the FM and the AF orders at $h=\pm 2$, respectively.
@@
$\mathcal{H}_{g}\simeq\frac{g\gamma}{\pi^{2}}\cos\sqrt{2\pi}\varphi_{s},$